DAILY SUMMARY REPORT SOL #14
Summary Title: We made it! Ready for re-entry!
Mission Status: Complete.
Sol Activity Summary: Our crew broke sim this morning, danced around outside in the snow and greeted the MDRS director as she arrived to the new barnacle behind the hab. We made a trip into town to drop off the empty food bins and made and also visited the Earth-based burger shack for some fresh food together. We returned to the hab to begin cleaning and packing as our mission is now complete!
Look Ahead Plan: We will re-enter Earth's atmosphere tomorrow afternoon!
Anomalies in work: Water pump issue, as reported in engineering report. MDRS director will assist us in investigation tomorrow morning in daylight. We have also reported that the hot water heater is not producing hot water and has actually not been functioning properly during our rotation. We will try to clean filter tomorrow and hand it over to the director for further action.
Weather: Cold, light snow, low wind.
Crew Physical Status: Crew has been taking vitamins and getting more rest and just generally taking care of itself.
EVA: No EVA performed today
Reports to be filed:
Support Requested: Mission Support, please inquire with Crew 161, Team Skoltech, as to whether or not their hot water functioned properly during their two-weeks.